# Fading ergodicity

### v1

@@ Abstract
Figure 4: Time evolution of the autocorrelation function $C(t)=\langle\hat{S}_{L}^{z}(t)\hat{S}_{L}^{z}(0)\rangle$ at $\alpha=0.86$ and different system sizes $L$, as a function of scaled time $\Gamma t$. Dashed line denotes the exponential decay $\propto e^{-\Gamma t}$, and horizontal dotted lines represent the long-time average $C_{\infty}$. Inset: difference $\Delta\mathcal{Q}_{\infty}$ between the long-time average of the observable $\mathcal{Q}(t)=\langle\hat{S}_{L}^{z}(t)\rangle$ after a quantum quench and the microcanonical ensemble prediction, vs $1/L$. See SM for details about the quench protocols.
Finally, we demonstrate the emergence of ergodicity from the dynamical perspective. In Fig. 4 we study the dynamics of two related quantities, the autocorrelation function $C(t)=\langle\hat{S}_{L}^{z}(t)\hat{S}_{L}^{z}(0)\rangle$ and the observable $\mathcal{Q}(t)=\langle\hat{S}_{L}^{z}(t)\rangle$ after a quantum quench from initial product states, while in SM we show the growth of the entanglement entropy towards the maximal value. The autocorrelation function $C(t)$ shown in the main panel of Fig. 4 exhibits an initial exponential decay, followed by an approach at long times to the steady-state value that vanishes with increasing the system size $L$. In the inset of Fig. 4, we compare the long-time value of $\mathcal{Q}(t)$ after a quantum quench to the microcanonical ensemble prediction, and we show that their difference vanishes as $L\to\infty$. In SM we also demonstrate that the temporal fluctuations above the long-time average vanish in the thermodynamic limit. Still, the finite-size scaling of temporal fluctuations exhibit softening that resembles the softening of the fluctuations exponent $\eta$ shown in Fig. 3(d). Since the temporal fluctuations are upper-bounded by the decay of the maximal off-diagonal matrix element dalessio_kafri_16, they represent an indicator to detect fingerprints of fading ergodicity based on quantum dynamics.
@@
Conclusions. Historically, the ETH ansatz for the matrix elements of observables was introduced by defining the smooth functions $O(\bar{E})$ and $f(\bar{E},\omega)$ as a refinement beyond the RMT behavior srednicki_99; dalessio_kafri_16. The physical significance of this refinement was to properly describe the energy dependence of matrix elements in physical systems that exhibit thermalization. Recent work has also explored refinements of ETH due to correlations between the matrix elements foini_kurchan_19, which give rise, among others, to nontrivial dynamics of four-point correlation functions foini_kurchan_19; chan_deluca_19; murthy_srednicki_19b; pappalardi_foini_22; Dymarsky2022. In this work we introduce a new modification of the RMT behavior that concerns the softening of the fluctuations of matrix elements. We refer to this phenomenon as fading ergodicity, which is characterized by the breakdown of the conventional ETH. Its physical significance is to allow for detecting boundaries of ergodicity. In case of a well-defined ergodicity breaking phase transition in the thermodynamic limit, fading ergodicity can be considered as a precursor of the transition point at which the ETH ceases to be valid.
@@ Acknowledgements.
We acknowledge discussions with J. Kurchan, P. Łydżba, M. Mierzejewski, A. Polkovnikov, M. Rigol and P. Sierant. We acknowledge support from the Slovenian Research and Innovation Agency (ARIS), Research core funding Grants No. P1-0044, N1-0273, J1-50005 and N1-0369. We gratefully acknowledge the High Performance Computing Research Infrastructure Eastern Region (HCP RIVR) consortium vega1 and European High Performance Computing Joint Undertaking (EuroHPC JU) vega2 for funding this research by providing computing resources of the HPC system Vega at the Institute of Information sciences vega3. Numerical calculations have been partly carried out using high-performance computing resources provided by the Wrocław Centre for Networking and Supercomputing. Supplemental Material: Fading ergodicity
Maksymilian Kliczkowski1,2,3,4, Rafał Świętek2,1,4, Miroslav Hopjan2 and Lev Vidmar2,1 1Department of Physics, Faculty of Mathematics and Physics, University of Ljubljana, SI-1000 Ljubljana, Slovenia 2Department of Theoretical Physics, J. Stefan Institute, SI-1000 Ljubljana, Slovenia 3Institute of Theoretical Physics, Faculty of Fundamental Problems of Technology,
Wrocław University of Science and Technology, 50-370 Wrocław, Poland
4These authors contributed equally: Maksymilian Kliczkowski, Rafał Świętek
S1 Matrix elements of observables
Figure S1: (a) A sketch of matrix elements $\langle n|O|m\rangle$ studied in this work. Main panel: The shaded region represents the off-diagonal matrix elements used to extract their $\omega$-dependence. Inset: diagonal matrix elements in the middle of the spectrum, around the mean energy $E_{\rm av}$. (b) Average gap ratio $r$ vs $\alpha$, for different system sizes $L$. Vertical dashed line denotes the transition point $\alpha_{c}$, extracted in Fig. 3(d) of the main text, while the horizontal solid and dotted lines show the ergodic and non-ergodic value respectively.
@@
Figure S2: (a)-(c) Coarse-grained off-diagonal matrix elements $\overline{|O_{nm}|^{2}}\rho$ vs $\omega$, where the density of states $\rho$ is obtained from a small energy window in the middle of the spectrum. Results are shown for the observables $\hat{S}^{z}_{L-1}\hat{S}^{z}_{L}$, $\hat{S}^{z}_{\rm dot}\hat{S}^{z}_{L}$, and $\hat{S}^{x}_{\rm dot}\hat{S}^{x}_{L}$, respectively, for different $L$. (d)-(f) Eigenstate-to-eigenstate fluctuations of the diagonal matrix elements, $\delta_{n}\equiv|O_{n+1}-O_{n}|$, vs $L$, for the same observables as shown in (a)-(c), respectively. We show the average values of $\delta_{n}$ in the main panels and the maximal outliers in the insets. Dashed lines are two-parameter fits to the function $a_{0}2^{-L/\eta}$.
S2 Quantum sun model
@@
$\hat{H}_{\rm dot}=\frac{1}{\sqrt{2^{N}+1}}R\;,$ (S2)
where $R=(A+A^{T})/\sqrt{2}$ and the matrix elements $A_{i,j}=\mathcal{N}(0,1)$ are sampled from a normal distribution with zero mean and unit variance. Each of the spins outside the dot is coupled to a single randomly selected spin within the dot with a strength of $\alpha^{u_{j}}$, where $\alpha$ is the parameter that drives the ergodicity breaking transition, and $u_{j}$ (with $j=1,2,...,L$) are drawn from a uniform distribution, $u_{j}\in[(j-1)-0.2,(j-1)+0.2]$, except for $j=1$ when $u_{1}=0$. The spins outside the dot are also subject to random magnetic fields $h_{j}$, which are drawn from a uniform distribution, $h_{j}\in[0.5,1.5]$. The diagonal and off-diagonal matrix elements of observables, discussed in Sec. S1, are collected with $N_{\rm samples}=2000$ for $11\geq L\geq 5$ and $N_{\rm samples}=874,89$ for $L=12,13$, respectively.
The analytical prediction for the ergodicity breaking transition point, based on the hybridization condition deroeck_huveneers_17, is $\tilde{\alpha}_{c}=1/\sqrt{2}\approx 0.707$. The exact numerical calculations of different ergodicity indicators predict the value of the transition point $\alpha_{c}$ that is close to $\tilde{\alpha}_{c}$. For the model parameters considered here the transition point is estimated to be in the interval $\tilde{\alpha}_{c}\lesssim\alpha_{c}\lesssim 0.75$ suntajs_hopjan_24. In particular, in Fig. 3(d) of the main text we obtained $\alpha_{c}\approx 0.74$ using the divergence of the fluctuation exponent $\eta$ as the transition indicator. In Fig. S1(b) we study the standard transition indicator, namely, the average nearest level spacing ratio $r$, shortly the average gap ratio. Introducing $r_{n}={\rm min}\{\delta E_{n},\delta E_{n-1}\}/{\rm max}\{\delta E_{n},\delta E_{n-1}\}$, where $\delta E_{n}=E_{n+1}-E_{n}$ is the level spacing between the levels $n$ and $n+1$, we obtain the average gap ratio $r$ by averaging $r_{n}$ over 500 eigenstates near the center of the spectrum and over different Hamiltonian realizations. Results in Fig. S1(b) suggest that the scale invariant point of $r$ vs $\alpha$, which denotes the ergodicity breaking transition point suntajs_hopjan_24, is quantitatively very close to $\alpha_{c}\approx 0.74$ predicted by the analysis of $\eta$ in Fig. 3(d).
S3 Choice of observables
@@
The results in Fig. S2 show how susceptible are the fluctuations of observable matrix elements to the proximity of the ergodicity breaking transition. The behavior of the matrix elements of $\hat{S}^{z}_{L-1}\hat{S}^{z}_{L}$ is similar to the behavior of $\hat{S}_{L}^{z}$ studied in the main text: the spectral weight of the off-diagonal matrix elements is accumulating at low $\omega$ [cf. Fig. S2(a)] and the fluctuations of the diagonal matrix elements gradually soften when $\alpha$ approaches the transition point [cf. Fig. S2(d)]. The common property of the observables $\hat{S}_{L}^{z}$ and $\hat{S}^{z}_{L-1}\hat{S}^{z}_{L}$ is that they measure properties of spins that are at large distance from the ergodic quantum dot, and hence they exhibit the weakest coupling in the system.
On the other hand, both observables $\hat{S}^{z}_{\rm dot}\hat{S}^{z}_{L}$ and $\hat{S}^{x}_{\rm dot}\hat{S}^{x}_{L}$ include the operator within the ergodic quantum dot. This property makes the proximity of the ergodicity breaking transition less apparent, as seen in Figs. S2(b)-S2(c) and S2(e)-S2(f). For the matrix elements of $\hat{S}^{z}_{\rm dot}\hat{S}^{z}_{L}$, the softening of fluctuations is mild but still noticeable [cf. Figs. S2(b) and S2(e)], while for $\hat{S}^{x}_{\rm dot}\hat{S}^{x}_{L}$, there is no signature of softening [cf. Figs. S2(c) and S2(f)]. The latter is a consequence of the nonergodic phase being localized in the computational basis suntajs_hopjan_24, which is the eigenstate basis of the $\hat{S}^{z}$ operator. For the observable $\hat{S}^{x}_{\rm dot}\hat{S}^{x}_{L}$, we find that the conventional ETH appears to be valid even on the non-ergodic side of the transition. Hence the matrix elements of the observable $\hat{S}^{x}_{\rm dot}\hat{S}^{x}_{L}$ are not expected to contain any special signature of the fading ergodicity regime.
S4 Distribution of matrix elements

### v2

@@ IV.2 Quantum dynamics
Figure 4: Quantum dynamics of the observable $\hat{S}_{L}^{z}$. (a) Difference $\Delta\mathcal{Q}_{\infty}$, see Eq. (12), between the microcanonical ensemble and the diagonal ensemble prediction after a quantum quench, vs $L$. Squares are results at $\alpha=0.74\approx\alpha_{c}$, at which the fluctuation exponent $\eta$ in Fig. 3(d) diverges, $\eta\to\infty$. Black circles are results for $\alpha=0.86$, which is studied in Fig. 2. (b) Time evolution of the autocorrelation function $C(t)$, see Eq. (13), at $\alpha=0.86$ and different system sizes $L$, as a function of scaled time $\Gamma t$. Dashed line denotes the exponential decay $\propto e^{-\Gamma t}$, and horizontal dotted lines represent the long-time average $C_{\infty}$.
Figure 4(a) shows the results for $\Delta\mathcal{Q}_{\infty}$ vs $L$ for different values of interaction $\alpha$. We observe that in the fading ergodicity regime at $\alpha>\alpha_{c}$, the differences $\Delta\mathcal{Q}_{\infty}$ appear to vanish in the thermodynamic limit $L\to\infty$. This suggests emergence of ergodicity for any finite value of the fluctuation exponent $\eta$. On the other hand, the infinite value of $\eta$, which we observe at $\alpha<\alpha_{c}$, see Fig. 3(d), does not ensure ergodicity. The breakdown of ergodicity is manifested in Fig. 4(a) for the results at $\alpha<\alpha_{c}$ (red dark circles above the squares), for which $\Delta\mathcal{Q}_{\infty}$ appears to approach a nonzero value in the thermodynamic limit.
Focusing on the dynamics at short and moderate times, we study a related quantity, the spin-spin autocorrelation function. We introduce it as
$C(t)=\langle\hat{S}_{L}^{z}(t)\hat{S}_{L}^{z}(0)\rangle_{\mu}\;,$ (13)
where the brackets $\langle\cdots\rangle_{\mu}$ denote both the quantum expectation value and the disorder average, see Eqs. (28) and (29) of Appendix E for the precise definition. The autocorrelation function $C(t)$ is shown in Fig. 4(b) in the ergodic regime at $\alpha=0.86$. It exhibits an exponential decay with the rate given by the Thouless energy $\Gamma$. At long times it approaches the steady-state value that vanishes with increasing the system size $L$, which is consistent with ergodic behavior.
Another important property of quantum quench dynamics is whether or not the expectation values of observables, such as $\mathcal{Q}(t)^{(\mu)}$, equilibrate after a sufficiently long time. This can be studied via the scaling of temporal fluctuations above the long-time average of expectation values. Our main question, however, is not whether the temporal fluctuations vanish in the thermodynamic limit (they do in the entire ergodic phase), but whether their finite-size scaling exhibits softening that resembles the softening of fluctuation exponents $\eta$ shown in Fig. 3(d). The affirmative answer to this question would imply the temporal fluctuations to be an indicator, based on quantum dynamics, for the detection of fading ergodicity. This expectation is based on their exact property (in the absence of degeneracies in the spectrum), i.e., they are upper-bounded by the decay of the maximal off-diagonal matrix element dalessio_kafri_16 .
Figure 5: Scaling of the long-time temporal fluctuations. (a) Scaling of the variance of temporal fluctuations $\sigma_{t}^{2}$, see Eq. (15), vs $L$, at different $\alpha$. We extract the exponent $\eta$ of the temporal fluctuations by fitting the function $b_{0}2^{-2L/\eta}$ to the numerical values (triangles). (b) Fluctuation exponents $\eta$ vs $\alpha$. The exponents of the temporal fluctuations are shown as triangles. Fluctuation exponents of the diagonal (squares) and the low-$\omega$ off-diagonal matrix elements (circles) are identical to those in Fig. 3(d).
To address this question, we define the variance of long-time temporal fluctuations for a given Hamiltonian realization $\hat{H}^{(\mu)}$,
$\sigma^{2}_{t,\mu}\equiv{\rm Var}_{t}\left\{\mathcal{Q}(t)^{(\mu)}\right\}\;,$ (14)
where the variance ${\rm Var}_{t}\{\cdots\}$ is calculated in a restricted time window after the Heisenberg time $t_{\rm H}$. Here, we select the time window $t\in[50t_{H},100t_{H}]$. The corresponding average variance over the Hamiltonian realizations is defined as
$\sigma^{2}_{t}={\rm Avr}_{\mu}\left\{\sigma^{2}_{t,\mu}\right\}\;.$ (15)
In Fig. 5(a) we show examples of $\sigma_{t}^{2}$ vs $L$ at different values of $\alpha$. We observe that in the entire ergodic regime, the decay of $\sigma_{t}^{2}$ is exponential with $L$, suggesting equilibration of the system.
Results in Fig. 5(a) motivate us to systematically study how the rate of the exponential decay of temporal fluctuations decreases when $\alpha$ approaches the ergodicity breaking transition. In Fig. 5(b) we compare the fluctuation exponents of temporal fluctuations with the fluctuations exponents of matrix elements fluctuations from Fig. 3(d). The obtained temporal fluctuation exponent is quantitatively smaller than the matrix element fluctuation exponent. However, in all cases, the fluctuation exponents increase when $\alpha$ approaches the transition point $\alpha_{c}$, suggesting that the long-time temporal fluctuations indeed carry information about the softening of the low-$\omega$ off-diagonal matrix elements. Hence, the breakdown of the conventional ETH in the fading ergodicity regime is also manifested in the scaling of the variance of temporal fluctuations with system size.
Figure 6: Time evolution of the entanglement entropy ${\cal S}_{L}(t)$, see Eq. (16), of a single (most weakly coupled) spin. Main panel: $\alpha=0.86>\alpha_{c}$ in the fading ergodicity regime. Inset: $\alpha=0.72<\alpha_{c}$ in the non-ergodic regime. Results are shown for various system sizes $L$. Solid lines represent the averaged entanglement entropy ${\cal S}_{L}(t)$, while the temporal fluctuations in different Hamiltonian realizations are denoted by fading colors.
IV.3 Growth of entanglement entropy
To complement the analysis, we study the time evolution of the entanglement entropy ${\cal S}_{L}(t)$. The goal is to show that in the fading ergodicity regime, despite the softening of the fluctuations of matrix elements, one still observes the growth of the entanglement entropy towards the maximal value. The von Neumann entanglement entropy ${\cal S}_{L}(t)$ is defined as
[…]
Results shown in Fig. 6 represent the entanglement growth in the ergodic regime (main panel) and in the non-ergodic regime (inset). We observe that in the ergodic regime, the entanglement entropy grows towards its maximum value, ${\cal S}(t\to\infty)\to\ln 2$. The approach towards maximal entanglement entropy is consistent with the ergodic behavior of the system (however, it is not a sufficient condition; see, e.g., Ref. vidmar_hackl_17 for a counter-example). On the other hand, the breakdown of ergodicity is clearly manifested in the inset of Fig. 6 at $\alpha<\alpha_{c}$, in which the entanglement entropy after a long time saturates to a value that is below the maximal value.
V Conclusions
@@
Historically, the ETH ansatz for the matrix elements of observables was introduced by defining the smooth functions $O(\bar{E})$ and $f(\bar{E},\omega)$ as a refinement beyond the RMT behavior srednicki_99 ; dalessio_kafri_16 . The physical significance of this refinement was to properly describe the energy dependence of matrix elements in physical systems that exhibit thermalization. Recent work has also explored refinements of ETH due to correlations between the matrix elements foini_kurchan_19 , which give rise, among others, to nontrivial dynamics of four-point correlation functions foini_kurchan_19 ; chan_deluca_19 ; murthy_srednicki_19b ; pappalardi_foini_22 ; Dymarsky2022 . In this work we introduce a new modification of the RMT framework that concerns the softening of the fluctuations of matrix elements. We refer to this phenomenon as fading ergodicity, which is characterized by the breakdown of the conventional ETH. Its physical significance is to allow for detecting boundaries of ergodicity. In case of a well-defined ergodicity breaking phase transition in the thermodynamic limit, fading ergodicity can be considered as a precursor of the transition point at which the ETH ceases to be valid.
@@ Acknowledgements.
We acknowledge discussions with J. Kurchan, P. Łydżba, M. Mierzejewski, A. Polkovnikov, M. Rigol and P. Sierant. We acknowledge support from the Slovenian Research and Innovation Agency (ARIS), Research core funding Grants No. P1-0044, N1-0273, J1-50005 and N1-0369. We gratefully acknowledge the High Performance Computing Research Infrastructure Eastern Region (HCP RIVR) consortium vega1 and European High Performance Computing Joint Undertaking (EuroHPC JU) vega2 for funding this research by providing computing resources of the HPC system Vega at the Institute of Information sciences vega3 . Numerical calculations have also been carried out using high-performance computing resources provided by the Wrocław Centre for Networking and Supercomputing.
Appendix A Matrix elements of observables
Figure 7: (a) A sketch of matrix elements $\langle n|O|m\rangle$ studied in this work. Main panel: The shaded region represents the off-diagonal matrix elements used to extract their $\omega$-dependence. Inset: diagonal matrix elements in the middle of the spectrum, around the mean energy $E_{\rm av}$. (b) Average gap ratio $r$ vs $\alpha$, for different system sizes $L$. Vertical dashed line denotes the transition point $\alpha_{c}$, extracted in Fig. 3(d) of the main text, while the horizontal solid and dotted lines show the ergodic and non-ergodic values, respectively.
@@ Appendix A Matrix elements of observables
Figure 8: (a)-(c) Coarse-grained off-diagonal matrix elements $\overline{|O_{nm}|^{2}}\rho$ vs $\omega$, where the density of states $\rho$ is obtained from a small energy window in the middle of the spectrum. Results are shown for the observables $\hat{S}^{z}_{L-1}\hat{S}^{z}_{L}$, $\hat{S}^{z}_{\rm dot}\hat{S}^{z}_{L}$, and $\hat{S}^{x}_{\rm dot}\hat{S}^{x}_{L}$, respectively, for different $L$. (d)-(f) Eigenstate-to-eigenstate fluctuations of the diagonal matrix elements, $\delta_{n}\equiv|O_{n+1}-O_{n}|$, vs $L$, for the same observables as shown in (a)-(c), respectively. We show the average values of $\delta_{n}$ in the main panels and the maximal outliers in the insets. Dashed lines are two-parameter fits to the function $a_{0}2^{-L/\eta}$.
Appendix B Quantum sun model
@@
$\hat{H}_{\rm dot}=\frac{1}{\sqrt{2^{N}+1}}R\;,$ (18)
where $R=(A+A^{T})/\sqrt{2}$ and the matrix elements $A_{i,j}=\mathcal{N}(0,1)$ are sampled from a normal distribution with zero mean and unit variance. Each of the spins outside the dot is coupled to a single randomly selected spin within the dot with a strength of $\alpha^{u_{j}}$, where $\alpha$ is the parameter that drives the ergodicity breaking transition, and $u_{j}$ (with $j=1,2,...,L$) are drawn from a uniform distribution, $u_{j}\in[(j-1)-0.2,(j-1)+0.2]$, except for $j=1$ when $u_{1}=0$. The spins outside the dot are also subject to random magnetic fields $h_{j}$, which are drawn from a uniform distribution, $h_{j}\in[0.5,1.5]$. The diagonal and off-diagonal matrix elements of observables, discussed in Appendix A, are collected using $N_{\rm samples}=2000$ for $11\geq L\geq 5$ and $N_{\rm samples}=874,89$ for $L=12,13$, respectively.
The analytical prediction for the ergodicity breaking transition point, based on the hybridization condition deroeck_huveneers_17 , is $\tilde{\alpha}_{c}=1/\sqrt{2}\approx 0.707$. The exact numerical calculations of different ergodicity indicators predict the value of the transition point $\alpha_{c}$ that is close to $\tilde{\alpha}_{c}$. For the model parameters considered here the transition point is estimated to be in the interval $\tilde{\alpha}_{c}\lesssim\alpha_{c}\lesssim 0.75$ suntajs_hopjan_24 . In particular, in Fig. 3(d) of the main text we obtained $\alpha_{c}\approx 0.74$ using the divergence of the fluctuation exponent $\eta$ as the transition indicator. In Fig. 7(b) we study the standard transition indicator, namely, the average nearest level spacing ratio $r$, shortly the average gap ratio. Introducing $r_{n}={\rm min}\{\delta E_{n},\delta E_{n-1}\}/{\rm max}\{\delta E_{n},\delta E_{n-1}\}$, where $\delta E_{n}=E_{n+1}-E_{n}$ is the level spacing between the levels $n$ and $n+1$, we obtain the average gap ratio $r$ by averaging $r_{n}$ over 500 eigenstates near the center of the spectrum and over different Hamiltonian realizations. Results in Fig. 7(b) suggest that the scale invariant point of $r$ vs $\alpha$, which denotes the ergodicity breaking transition point suntajs_hopjan_24 , is quantitatively very close to $\alpha_{c}\approx 0.74$ predicted by the analysis of $\eta$ in Fig. 3(d).
Appendix C Choice of observables
@@
The results in Fig. 8 show how susceptible are the fluctuations of observable matrix elements to the proximity of the ergodicity breaking transition. The behavior of the matrix elements of $\hat{S}^{z}_{L-1}\hat{S}^{z}_{L}$ is similar to the behavior of $\hat{S}_{L}^{z}$ studied in the main text: the spectral weight of the off-diagonal matrix elements is accumulating at low $\omega$ [cf. Fig. 8(a)] and the fluctuations of the diagonal matrix elements gradually soften when $\alpha$ approaches the transition point [cf. Fig. 8(d)]. The common property of the observables $\hat{S}_{L}^{z}$ and $\hat{S}^{z}_{L-1}\hat{S}^{z}_{L}$ is that they measure properties of spins that are at large distance from the ergodic quantum dot, and hence they exhibit the weakest coupling in the system.
On the other hand, both observables $\hat{S}^{z}_{\rm dot}\hat{S}^{z}_{L}$ and $\hat{S}^{x}_{\rm dot}\hat{S}^{x}_{L}$ include the operator within the ergodic quantum dot. This property makes the proximity of the ergodicity breaking transition less apparent, as seen in Figs. 8(b)-8(c) and 8(e)-8(f). For the matrix elements of $\hat{S}^{z}_{\rm dot}\hat{S}^{z}_{L}$, the softening of fluctuations is mild but still noticeable [cf. Figs. 8(b) and 8(e)], while for $\hat{S}^{x}_{\rm dot}\hat{S}^{x}_{L}$, there is no signature of softening [cf. Figs. 8(c) and 8(f)]. The latter is a consequence of the nonergodic phase being localized in the computational basis suntajs_hopjan_24 , which is the eigenstate basis of the $\hat{S}^{z}$ operator. For the observable $\hat{S}^{x}_{\rm dot}\hat{S}^{x}_{L}$, we find that the conventional ETH appears to be valid even on the non-ergodic side of the transition. Hence the matrix elements of the observable $\hat{S}^{x}_{\rm dot}\hat{S}^{x}_{L}$ are not expected to contain any special signatures of the fading ergodicity regime.
Appendix D Distribution of matrix elements